If 3.58 g CuNO3is dissolved in water to make a 0.670 M solution, what is the volume of the solution?
1
Expert's answer
2019-08-06T02:04:04-0400

Using equation for molar concentration "c=\frac{n}{V}"

, where c - molar concentration, n - amount of compound in moles, V - volume of solution

we can determine volume of solution:

"V(CuNO_3)=\frac{n(CuNO_3)}{c(CuNO_3)}"

The amount n(CuNO3) can be found from the equation

"n(CuNO_3)=\frac{m(CuNO_3)}{M(CuNO_3)}"

,where m - mass of CuNO3 and M(CuNO3) is molar mass of CuNO3.

Thus "V(CuNO_3)=\frac{n(CuNO_3)}{c(CuNO_3)}=\frac{m(CuNO_3)}{M(CuNO_3) \cdot c(CuNO_3)}=\frac{3.58g}{126\frac{g}{mol} \cdot 0.670 \frac{mol}{L}}=0.042L=42mL"


By the way, copper (I) nitrate CuNO3 doesn't exist in pure form without ligands, thus the problem in the given wording doesn't make any practical sense.